OpenStudy (anonymous):

R=(Vi^2)/g*(sin(2theta)) solve for theta?

OpenStudy (barrycarter):

Hint: arcsin

OpenStudy (anonymous):

You must rewrite the expression \[\theta=\frac{1}{2}\sin^{-1} \left(\frac{ gR }{ v_i^2 }\right)\]